How do you factor this expression 4x^2+12x+9?
Answer by kiru_khandelwal(79)

4x^2 + 12x + 9 = 0
=> 4x^2 + 6x + 6x + 9 =0
=> 2x(2x+3) + 3(2x+3) =0
=>(2x+3)(2x+3) = 0
=> 2x+3 = 0
=> x = -3/2